# Mitigating Image Captioning Hallucinations in Vision-Language Models

### v1

@@ I Introduction
The root cause of hallucinations often lies in distribution shifts between pretraining data and real-world test samples. Pretraining on noisy datasets introduces biases and reliance on unimodal priors, leading models to generate hallucinated content when presented with unseen or domain-specific data. Addressing this challenge requires methods that adapt models to new data distributions while ensuring computational efficiency.
Existing methods to mitigate hallucinations, including retraining or fine-tuning, ensemble approaches, and logit manipulation, face limitations in scalability, computational efficiency, and generalizability (see Section II).
@@
    A Parameter-Efficient RL Approach for VLMs: Our approach updates only the learnable parameters in layer normalization [3], accounting for approximately 0.003% of the model’s total parameters. This is significantly more efficient than modifying entire VLMs or their cross-modal projection layers that have nearly 100 times more parameters.
@@
The remainder of this paper is organized as follows: Section II reviews related work on hallucination mitigation in VLMs. Section III details our proposed methodology, including the TTA framework, the RL setup, and the customized CLIP-based hallucination evaluation model. Section IV describes the datasets used for experiments, while Section V presents experimental results and offers ablation studies to further highlight the contributions of individual components. Finally, Section VI concludes the paper and discusses potential future directions.
@@ III-B Reward Calculation
where $\beta$ is a weighting factor. This dual-reward system ensures that the VLM generates captions that are more semantically aligned with the input image with less hallucination, making it a crucial component of the RL framework.
@@ III-C Hallucination Evaluation Model
The hallucination evaluation model is built upon the frozen CLIP architecture, enhanced with a single learnable query token ${Q}$. This token is updated iteratively through $N$-stage cross-attention [21] layers to extract alignment and semantic consistency features between image and text inputs. The final updated query token features (${Q}_{{N}}$) serve as input to a classifier, composed of stacked fully connected layers, which predicts whether an image-text pair is non-hallucinated (positive) or hallucinated (negative). The classifier’s output, a logit score referred to as the NHP score, serves as a crucial reward in the RL framework.
@@
where ${Q}_{i-1}$ is the query token from the previous stage, and ${K}_{i},{V}_{i}$ represent the keys and values derived from the concatenated image ($v$) and text ($y$) features at stage $i$.
The CLS token (classification token representing the image’s global features) from the image encoder and the EOS token (End-of-Sequence token) from the text encoder are used to compute the SAS score via their cosine similarity, providing a quantitative measure of alignment between the image and the caption.
@@ V Results and Ablation Studies
As shown in Table II, the CLIPPrompts + Triplet model comprises only 5.8M learnable parameters, emphasizing its lightweight design. Before adopting it in our RL framework, we explored two variants. In CLIPstacked, CLS and EOS token features are combined and passed into stacked fully connected layers for classification. In contrast, CLIPPrompts introduces a learnable token that interacts with the frozen encoders, producing features then fed into stacked fully connected layers. Although both variants outperform the LLaVA models, neither surpasses our final CLIPPrompts + Triplet model, which builds upon CLIPPrompts by incorporating a triplet loss (see Section III-C) to further refine the representation space. As shown in Table V, CLIPPrompts + Triplet achieves the highest F1 scores for Object (86.5%), Attribute (80.7%), and Relation (75.5%), surpassing the other models. These results highlight the synergistic effect of combining learnable prompts with triplet loss, enabling the CLIP-based model to deliver superior accuracy and robustness in hallucination evaluation while maintaining efficiency.
@@ VI Conclusion and Future Directions
This work tackles the critical challenge of mitigating hallucinations in VLMs by introducing a novel TTA framework using RL. By updating only the lightweight layer normalization gamma parameters and incorporating a customized CLIP-based hallucination evaluation model, our approach effectively reduces object hallucinations during inference. Experiments demonstrate substantial performance improvements across multiple metrics on SOTA VLMs, showcasing the framework’s robustness and efficiency.
In future work, we aim to extend this framework to attribute and relationship hallucinations, leveraging the capabilities of the customized CLIP model for a more comprehensive solution. Additionally, we plan to explore its application in discriminative tasks and adversarial training setups, where the VLM acts as a generative model and the evaluation model serves as a discriminator. These advancements may further enhance hallucination mitigation in VLMs, broadening their real-world applicability.

### v2

@@ I Introduction
The root cause of hallucinations often lies in distribution shifts between pretraining data and real-world test samples. Pretraining on noisy datasets introduces biases and reliance on unimodal priors, leading models to generate hallucinated content when presented with unseen or domain-specific data. Addressing this challenge requires methods that adapt models to new data distributions while ensuring computational efficiency. Existing methods to mitigate hallucinations, including retraining or fine-tuning, ensemble approaches, and logit manipulation, face limitations in scalability, computational efficiency, and generalizability (see Section II).
@@
    A Parameter-Efficient RL Approach for VLMs: Our approach updates only the learnable parameters in layer normalization [3], accounting for approximately 0.003% of the model’s total parameters. This is significantly more efficient than modifying entire VLMs or cross-modal projection layers that have nearly 100 times more parameters.
@@
The remainder of this paper is organized as follows: Section II reviews related work on hallucination mitigation in VLMs. Section III details our proposed methodology, including the TTA framework, the RL setup, and the customized CLIP-based hallucination evaluation model. Section IV describes the datasets, while Section V presents experimental results and offers ablation studies to further highlight the contributions of individual components. Finally, Section VI concludes the paper and discusses potential future directions.
@@ III-B Reward Calculation
where $\beta$ is a weighting factor, empirically set to 1. This dual-reward system ensures that the VLM generates captions that are more semantically aligned with the input image with less hallucination, making it a crucial component of the RL framework.
@@ III-C Hallucination Evaluation Model
The hallucination evaluation model builds on CLIP [21], a widely used standard for image-text alignment with strong zero-shot generalization. We introduce a single learnable query token ${Q}$ atop the frozen CLIP architecture. This token is updated iteratively through $N$-stage cross-attention [22] layers to extract alignment and semantic consistency features between image and text inputs. The final updated query token features (${Q}_{{N}}$) serve as input to a classifier, composed of stacked fully connected layers, which predicts whether an image-text pair is non-hallucinated (positive) or hallucinated (negative). The classifier’s output, a logit score referred to as the NHP score, serves as a crucial reward in the RL framework.
@@
where ${Q}_{i-1}$ is the query token from the previous stage, and ${K}_{i},{V}_{i}$ represent the keys and values derived from the concatenated image ($v$) and text ($y$) features at stage $i$. The CLS token (classification token representing the image’s global features) from the image encoder and the EOS token (End-of-Sequence token) from the text encoder are used to compute the SAS score via their cosine similarity, providing a quantitative measure of alignment between the image and the caption.
@@ V Results and Ablation Studies
As shown in Table II, the CLIPPrompts + Triplet model comprises only 5.8M learnable parameters, emphasizing its lightweight design. Before adopting it in our RL framework, we explored two variants. In CLIPstacked, CLS and EOS token features are combined and passed into stacked fully connected layers for classification. In contrast, CLIPPrompts introduces a learnable token that interacts with the frozen encoders, producing features then fed into stacked fully connected layers. Although both variants outperform the LLaVA models, neither surpasses our final CLIPPrompts + Triplet model, which builds upon CLIPPrompts by incorporating a triplet loss (see Section III-C) to further refine the representation space. As shown in Table V, CLIPPrompts + Triplet achieves the highest F1 scores for Object (86.5%), Attribute (80.7%), and Relation (75.5%), surpassing the other models. It highlights the synergistic effect of combining learnable prompts with triplet loss, enabling the CLIP-based model to deliver superior accuracy and robustness in hallucination evaluation while maintaining efficiency.
@@ VI Conclusion and Future Directions
This work tackles the critical challenge of mitigating hallucinations in VLMs by introducing a novel TTA framework using RL. By updating only the lightweight layer normalization gamma parameters and incorporating a customized CLIP-based hallucination evaluation model, our approach effectively reduces object hallucinations during inference. Experiments demonstrate substantial performance improvements across multiple metrics on SOTA VLMs, showcasing the framework’s robustness and efficiency. In future work, we aim to extend this framework to attribute and relationship hallucinations, leveraging the capabilities of the customized CLIP model for a more comprehensive solution. Additionally, we plan to explore its application in discriminative tasks and adversarial training setups, where the VLM acts as a generative model and the evaluation model serves as a discriminator. These advancements may further enhance hallucination mitigation in VLMs, broadening their real-world applicability.